# On the approximation of finite-time Lyapunov exponents for the stochastic Burgers equation

### v1

@@ 5 Bifurcations for one-dimensional kernel
To simplify the results we assume throughout the section that $\text{dim}(\mathcal{N})=1$. Now according to Assumption 2.3
@@ 5.2 Case $1\gg\sigma\approx-\nu>0$
and transfer it to the SDE by LDP. As we can rely on the deterministic amplitude equation (neglect the small noise term), we do not need the support theorem and could prove a result for instability or stability (depending on the sign of $\nu$) also with probability almost one, depending on the initial condition. We refrain from giving more details here.
@@
For $\nu=0$ we could use either the ergodic theorem [BlNe:23, Lemma 4.4] if $a_{0}$ is distributed according to the the invariant measure of the AE to show that $\lambda_{T}^{a}<0$ with high probability, or for general $a_{0}$ we can rely on an argument of [BB:25, Lemma 4.2].
In both cases we can carry over this stability to the SPDE using Corollary 4.7.
For the case $\nu\not=0$ we need an approximation result that relates results of $\nu=0$ to the non-zero but small case, which we will not state here.
@@ 6.1 Burgers equation
$W(t)=\sum\limits_{k=1}^{\infty}\alpha_{k}\beta_{k}(t)e_{k},$
where $(\alpha_{k})_{k\geq 1}$ is a sequence of bounded real numbers and $(e_{k})_{k\geq 1}$ is the ONB specified above. The covariance operator $Q$ is Fourier diagonal and satisfies $Qe_{k}=\alpha^{2}_{k}e_{k}$, which implies that $W_{c}(t)=\alpha_{1}\beta_{1}(t)e_{1}$. Plugging this in (2.3) and (2.4) verifies Assumption 2.4. We also note the $A_{s}^{-1}$ is a trace class operator, which shows that $W_{s}$ is an $X$-valued process. Moreover, the assumption on the OU-process in Assumption 2.5 is well-known in the literature [DaPZ:92] and straightforward to check.
@@
and the bifurcation analysis made Section 5 applies in this case.
@@ 6.2 Kuramoto-Sivashinsky
The operator $A=-\partial_{x}^{4}$ satisfies Assumption 2.2 as it is Fourier diagonal $\mathcal{N}=\text{span}\{1\}$. The spaces are $X=L^{2}([0,2\pi])$ and $Y=H^{5/4}_{\text{per}}([0,2\pi])$.
One can readily check that for $B(u,v)=\partial_{x}u\cdot\partial_{x}v$ satisfies Assumption 2.3 and with a Fourier-diagonal Wiener process $W$ as in the example above both Assumptions 2.4 and 2.5 are easy to check in a similar way as for Burgers equation.
Unfortunately, $\mathcal{F}_{c}(a)=0$ and thus the amplitude equation is only the linear equation
${\textnormal{d}}a=\frac{\nu}{\varepsilon^{2}}a\penalty 10000\ {\textnormal{d}}T+\frac{\sigma}{\varepsilon^{2}}\alpha_{1}\penalty 10000\ {\textnormal{d}}\beta_{\varepsilon}(T)$ (6.4)
@@ 6.3 Rayleigh-Bénard
This model consists of a Navier-Stokes equation coupled to a heat equation, it does not fit into our setting, as the the operator $A$ is not self-adjoint but it has a compact resolvent. We have a basis of eigenfuctions but this is not orthonormal, and we have to modify our result, for example $P_{c}$ is no longer an orthogonal projection with norm $1$, but this should be fine.
Moreover, $\mathcal{N}$ might not be one-dimansional and thus the bifurcation analysis presented in Section 5 does not apply in this case, as the FTLE for higher dimensional SDEs are in general not easy to compute.
Another example is a surface growth model in [BlMo:09] with periodic boundary conditions where the kernel is two-dimensional, here in some cases we might exploit symmetries like polar coordinates to reduce the AE from 2D to 1D.

### v2

@@ 5 Example: Bifurcations for one-dimensional kernel
To simplify the presentation of the results we assume throughout the section that $\text{dim}(\mathcal{N})=1$. Now according to Assumption 2.3
@@ 5.2 Case $1\gg\sigma\approx-\nu>0$
and transfer it to the SDE by LDP. As we can rely on the deterministic amplitude equation (neglecting the small noise term), we do not need the support theorem and could prove a result for instability or stability (depending on the sign of $\nu$) also with probability almost one, depending on the initial condition. We refrain from giving more details here.
@@
For $\nu=0$ we could use either Birkhoff’s ergodic theorem [BlNe:23, Lemma 4.4] if the initial datum $a_{0}(\omega)$ is distributed according to the the invariant measure of the AE to show that $\lambda_{T}^{a}<0$ with high probability. For general initial data $a_{0}$ we can rely on [BB:25, Lemma 4.2] to show that the solution of the AE does not spend too much time in zero from which we can infer again that $\lambda_{T}^{a}<0$ with high probability. In both cases we can carry over this stability to the SPDE using Corollary 4.7.
For the case $\nu\not=0$ we need an approximation result similar to Theorem 3.9 that relates results of $\nu=0$ to $\nu\neq 0$ but small, which we will not state here.
@@ 6.1 Burgers equation
$W(t)=\sum\limits_{k=1}^{\infty}\alpha_{k}\beta_{k}(t)e_{k},$ (6.3)
where $(\alpha_{k})_{k\geq 1}$ is a sequence of bounded real numbers and $(e_{k})_{k\geq 1}$ is the ONB specified above. The covariance operator $Q$ is Fourier diagonal and satisfies $Qe_{k}=\alpha^{2}_{k}e_{k}$, which implies that $W_{c}(t)=\alpha_{1}\beta_{1}(t)e_{1}$. Plugging this in (2.3) and (2.4) verifies Assumption 2.4. We also note the $A_{s}^{-1}$ is a trace class operator, which shows that $W_{s}$ is an $X$-valued process. Moreover, Assumption 2.5 is well-known in the literature [DaPZ:92] and straightforward to check.
@@
for which the bifurcation analysis made Section 5 applies.
@@ 6.2 Kuramoto-Sivashinsky
The operator $A=-\partial_{x}^{4}$ satisfies Assumption 2.2 as it is Fourier diagonal with $\mathcal{N}=\text{span}\{1\}$. The spaces are $X=L^{2}([0,2\pi])$ and $Y=H^{5/4}_{\text{per}}([0,2\pi])$. One can readily check that $B(u,v)=\partial_{x}u\cdot\partial_{x}v$ satisfies Assumption 2.3. Moreover, both Assumptions 2.4 and 2.5 are easy to check in a similar way as for Burgers equation with the same choice of noise as in (6.3). Unfortunately, $\mathcal{F}_{c}(a)=0$ leading to the linear amplitude equation
${\textnormal{d}}a=\frac{\nu}{\varepsilon^{2}}a\penalty 10000\ {\textnormal{d}}T+\frac{\sigma}{\varepsilon^{2}}\alpha_{1}\penalty 10000\ {\textnormal{d}}\beta_{\varepsilon}(T).$ (6.5)
@@ 6.3 Rayleigh-Bénard
This model consists of a Navier-Stokes equation coupled to a heat equation, it does not fit into our setting, as the the operator $A$ is not self-adjoint but it has a compact resolvent. We have a basis of eigenfunctions but this is not orthonormal, in particular $P_{c}$ is no longer an orthogonal projection with norm one. Moreover, $\mathcal{N}$ might not be one-dimensional and thus the bifurcation analysis presented in Section 5 does not apply in this case, as the FTLEs for higher dimensional SDEs are in general not easy to compute.